Paul and Maureen, two graduates at University, produce customized personal computer systems at night. They put together generic personal computers, which have various special features for students, and local businesses. They have recently received their biggest order to date from the EM department for 36 customized personal computers. Paul and Maureen assembled the first unit as a trial and found that it took them 18 hours of direct labor. To determine if they can fill the order in the time allotted, they want to apply the learning curve effect to determine how much time the 9th, 18th, and 36th units will require to assemble. Based on past experience, they believe their learning curve is 80%
Answer #1

learning curve = 80%, So b =log 0.8/log 2 => -0.322

Learning curve equation Y=axb

where Y = Total hours to produce nth unit

a = hours reqd for first unit

x = nth unit

b = learning curve

  1. 9th unit = (18) (9)^-0.322 => 8.87 hours
  2. 18th unit = (18) (18)^-0.322= 7.10 hours
  3. 36th unit = (18) (36)^-0.322 = 5.67 hours